# Reference-based Metrics Disprove Themselves in Question Generation

### v2

@@ 4.2 Results
BART-base 12.97 2.78 41.25 2.35
GPT-3.5 (zero-shot) 7.94 4.25 73.60 4.20
@@
4.3 Analysis
NACo vs. Reference-based Metrics: Tbl. 3 indicates that reference-based metrics rate BART-base questions slightly lower than AnnoQ (by $0.67\%$ according to BLEU), whereas NACo shows a much larger gap ($42.18\%$). Upon manually reviewing the questions generated by BART-base, we noticed a considerable number of them were not actual questions but rather statements using similar wording to the reference question RefQ. This observation is validated by our human evaluators, detailed in A.4.2. Fig. 3 provides a case study where BERTScore, the reference-based metric most aligned with human judgment Mohammadshahi et al. (2023), favored BART-base generation over the human annotated question, even though the former was not formatted as a question. This incompetency of reference-based metric can be explained by the fact that BART, when finetuned on the HotpotQA training set, can identify words that will be used in the reference RefQ, but fail to form a coherent and answerable question. NACo, emphasizing essential criteria of a question, assigns a score of 00 to the BART-base output while giving a high score for AnnoQ ($88.89$).
@@ 4.3 Analysis
NACo (CoT-QA) vs. LLM Direct Evaluation: Large language models (LLMs) are increasingly utilized as proxies for human evaluators. We examine the effectiveness of CoT-QA used by NACo, against the conventional direct use of LLMs for evaluation. In direct evaluation, the LLMs receive the same instruction/prompt as our human evaluators. Fig. 4 presents the Pearson correlation coefficients, comparing the performance of CoT-QA (NACo) with direct evaluation across individual criteria and overall question quality. The results indicate a higher alignment with human judgment when employing CoT-QA for each respective LLM. Notably, adopting CoT-QA instead of direct evaluation significantly boosts the performance of Claude-Haiku, improving the alignment with human judgment of overall question quality from $0.46$ to $0.73$. This improvement is comparable to the performance achieved using GPT4o ($0.69$ for direct evaluation, $0.80$ in the CoTQA setting), while being $12$ times more cost effective.
@@ A.4.2 Human Evaluation Results
BART-base 0.73 0.60 0.46 2.35
GPT-3.5 1.69 1.40 0.85 4.20
RefQ 1.48 1.46 1.35 4.77
AnnoQ 1.63 1.69 1.46 5.02
@@ A.8 Additional results for reference-based metrics
Tbl. 5 and 6 provides a more detailed version of 1 and 3.
Tbl. 7 illustrates the application of our study that disproves reference-based metrics and the proposed metric, NACo, in an educational setting using the TedEd-QA dataset. It can be seen that our observations regarding the failure of reference-based metrics and the effectiveness of NACo also holds for this dataset. Specifically, Refq, AnnoQ, and AnnoQ-clue-RefQ have significant gap when reference-based metrics are used to score them. NACo is able to score all these three human-validated candidates with similar scores and no worse than any machine-generated candidates.
@@ B B-RT R-L BSc Q-B QA-S R-Q NACo
BART-base 12.97 -0.79 34.38 87.88 27.67 -0.28 2.78 41.25 2.35
GPT-3.5 (zero-shot) 7.94 -0.95 26.05 87.38 17.00 -0.28 4.25 73.60 4.20
GPT-3.5 (few-shot) 10.53 -0.87 27.51 87.76 18.67 -0.27 4.21 71.09 -
BART-clue-RefQ 31.90 -0.10 58.77 92.61 51.98 -0.28 3.31 63.17 -
@@
RefQ 100.00 1.00 100.00 100.00 100.00 -0.27 4.21 74.33 4.77
AnnoQ 13.64 -0.70 30.94 89.22 26.95 -0.27 4.32 83.43 5.02
AnnoQ-clue-RefQ 35.59 -0.06 52.95 92.15 51.76 -0.27 4.44 81.60 -

### v3

@@ 4.3 Analysis
BART-base 14.57 2.90 42.65 2.80
GPT-3.5 (zero-shot) 9.46 4.18 74.99 4.60
@@
NACo vs. Reference-based Metrics: Tbl. 3 indicates that reference-based metrics rate BART-base questions slightly lower than AnnoQ (by $0.23\%$ according to BLEU), whereas NACo shows a much larger gap ($42.32\%$). Upon manually reviewing the questions generated by BART-base, we noticed a considerable number of them were not actual questions but rather statements using similar wording to the reference question RefQ. This observation is validated by our human evaluators, detailed in A.4.2. Fig. 3 provides a case study where BERTScore, the reference-based metric most aligned with human judgment Mohammadshahi et al. (2023), favored BART-base generation over the human annotated question, even though the former was not formatted as a question. This incompetency of reference-based metric can be explained by the fact that BART, when finetuned on the HotpotQA training set, can identify words that will be used in the reference RefQ, but fail to form a coherent and answerable question. NACo, emphasizing essential criteria of a question, assigns a score of 00 to the BART-base output while giving a high score for AnnoQ ($88.89$).
@@
Human Preference Study: As our human evaluation study assesses candidate questions based on the criteria that NACo measures, we conduct a human preference study to ensure fair comparisons between NACo and existing metrics. In this study, we further compare NACo with the top two baselines: RQUGE and BERTScore. We sample 20 pairs of questions where NACo and each baseline disagree in their scoring. Three human evaluators make their preferences for each pair, achieving a Cohen’s Kappa of 0.87, indicating strong agreement. Using human preference as the reference, NACo wins against RQUGE in 15 out of 20 cases (75%) and BERTScore in 12 out of 20 cases (60%).
NACo (CoT-QA) vs. LLM Direct Evaluation: Large language models (LLMs) are increasingly utilized as proxies for human evaluators. Previous studies have suggested that when receiving CoT instructions typically given to human evaluators, LLMs can assess generated texts in a way that are highly aligned with human judgement Liu et al. (2023). We refer to this approach of using LLM evaluators as Direct Evaluation (DirectEval). We examine the effectiveness of CoT-QA, used by NACo, against DirectEval, which provides the LLMs the same human evaluation instructions in Appx. A.4.1. Fig. 4 presents the Pearson correlation coefficients, comparing the performance of CoT-QA (NACo) with DirectEval across individual criteria and overall question quality. The results indicate a higher alignment with human judgment when employing CoT-QA for each respective LLM. Notably, adopting CoT-QA instead of DirectEval significantly boosts the performance of Claude-Haiku, improving the alignment with human judgment of overall question quality from $0.42$ to $0.71$. This improvement is comparable to the performance achieved using GPT4o ($0.72$ in DirectEval setting, $0.80$ in CoTQA setting), while being $12$ times more cost effective.
We also investigate whether DirectEval and NACo will benefit from the addition of a reference question during evaluation. For DirectEval, we provide the reference question at the end of the instruction. For NACo, we use the reference question’s complexity (obtained from CoT-QA) as the expected complexity. In our experiment, we use the RefQ group as the reference, and evaluate the other three groups of candidate questions (BART-base, GPT3.5, and AnnoQ). Table 4 details the correlation between human judgement and the use of LLM evaluators with and without references.
Method GPT3.5 Claude-Haiku
DirectEval 0.63 0.48
DirectEval + RefQ 0.74 0.52
NACo 0.72 0.74
NACo + RefQ 0.69 0.75
Table 4: Pearson correlation between human assessments and LLM evaluators, with and without references (RefQ).
It can be seen that while references benefit DirectEval to some extent, they do not consistently improve it to the level of NACo. Specifically, DirectEval $+$ RefQ with GPT3.5 shows on-par performance with NACo (0.74 v.s. 0.72), while DirectEval $+$ RefQ with Claude-Haiku underperforms NACo by a large margin. Furthermore, adding references on top of NACo does not further improve its performance. These results suggest that a single reference does not provide orthogonal benefits to NACo, which aligns with our findings regarding the limitations of reference-based metrics. NACo, as a reference-free metric, provides comprehensive and robust QG evaluations, without suffering from the bias of a single reference and the expensive reference collection process.
@@ A.4.2 Human Evaluation Results
BART-base 1.10 0.74 0.97 2.80
GPT-3.5 1.92 1.62 1.06 4.60
RefQ 1.70 1.68 1.75 5.14
AnnoQ 1.82 1.83 1.80 5.45
@@ A.8 Additional results for reference-based metrics
Tbl. 6 and 7 provides a more detailed version of 1 and 3.
Tbl. 8 illustrates the application of our study that disproves reference-based metrics and the proposed metric, NACo, in an educational setting using the TedEd-QA dataset. It can be seen that our observations regarding the failure of reference-based metrics and the effectiveness of NACo also holds for this dataset. Specifically, Refq, AnnoQ, and AnnoQ-clue-RefQ have significant gap when reference-based metrics are used to score them. NACo is able to score all these three human-validated candidates with similar scores and no worse than any machine-generated candidates.
@@ B B-RT R-L BSc Q-B QA-S R-Q NACo
BART-base 14.57 -0.75 34.87 87.83 28.58 -0.28 2.90 42.26 2.80
GPT-3.5 (zero-shot) 9.46 -0.92 26.95 87.48 16.87 -0.28 4.18 74.99 4.60
GPT-3.5 (few-shot) 9.55 -0.86 27.48 87.71 17.33 -0.27 4.33 77.41 -
BART-clue-RefQ 34.29 -0.07 61.40 92.86 53.94 -0.28 3.51 62.40 -
@@
RefQ 100.00 1.00 100.00 100.00 100.00 -0.28 4.12 75.32 5.14
AnnoQ 14.80 -0.59 34.21 89.65 28.77 -0.27 4.21 84.97 5.45
AnnoQ-clue-RefQ 32.56 -0.05 53.21 93.12 52.85 -0.27 4.26 79.16 -